November 1996  ICTCM IX (Reno)  15-1

Recall the variety of textbook problems. Some contained rather diverse sets of parameters. Are all realistic? An answer to this question requires some knowledge about the physics of falling objects and skydiving practices and procedures. This information is not difficult to find and is easily understandable – by students (and faculty)!

We need to be looking both for realistic parameter values and for indications that the model is correct. \underline{Definitions}:

• snatch force is acceleration at instant parachute deployment begins

• opening shock is force produced while parachute deploys

• jerk is rate of change of acceleration (opening shock during deployment)